1. Suppose a population of organisms is in Hardy–Weinberg equilibrium. In the population, a gene has two alleles, B and b. If the frequency of B is 0.72 and the frequency of b is 0.28, what are the expected genotype frequencies of BB, Bb, and bb? Calculate each frequency to two decimal places.
Answer #1

Answer:

p2 + 2pq + q2 = 1

p= 0.72 (B)

q= 0.28 (b)

Therefore,

(0.72)2 + 2 (0.72) (0.28) + (0.28)2

ƒ(BB)= p2 (0.72)2 = 0.52

ƒ(Bb)= 2pq= 2×(0.72)×(0.28) = 0.40

ƒ(bb)= q2 = (0.28)2 = 0.08

BB= 52%; Bb= 40%; bb= 8%